OpenStudy (anonymous):

three side a,b,and c of a triangle ABC satisfy the ratio 6b+c:3b+a:2c-a=4:2:1. If the area of triangle is 8sqrt{14},find the values of a,b and c

OpenStudy (wolf1728):

Is this the question ? 

OpenStudy (anonymous):

yes

OpenStudy (wolf1728):

Okay just thought I'd make really sure before I make any drawings, calculations, etc.

OpenStudy (anonymous):

'kay

OpenStudy (wolf1728):

A) 6b + c = 8c - 4a B) 3b + a = 4c -2a C) 6b + c = 6b + 2a Changing equation C c = 2a Equation B 3b + a = 4*2a -2a 3b = 5a Equation A 6b + c = 8c -4*(c/2) 6b -7c = -2c 6b = 5c -2a + c =0 -5a + 3b=0 6b -5c =0 I cannot solve these equations. Perhaps I didn't set up these equations properly.

OpenStudy (anonymous):

are u familiar with heron's formula for area of a triangle?

OpenStudy (wolf1728):

Yes - set up a semi-perimeter, etc

OpenStudy (anonymous):

right, suppose that sides are x, 2x and 4x then using heron's formula you can evaluate x

OpenStudy (wolf1728):

okay - good thought

OpenStudy (anonymous):

and then set up your equations like this 2c-a=x , 3b+a=2x and 6b+c=4x

OpenStudy (wolf1728):

Hey I just thought of something - the 3 sides cannot be in the ratio of 4 , 2 and 1 The longest side (4) must be smaller than the sum of the other 2 sides. It isn;t. ARRGGGHHHHHHH !!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

good point :-))

OpenStudy (anonymous):

x+2x < 4x

OpenStudy (wolf1728):

I was wondering why I couldn't solve it. Now I know.

OpenStudy (raden):

according what you got, @wolf1728 a = a c = 2a b = 5a/3 perimeter = a+b+c= a + 5a/3 + 2a = 14a/3 s =1/2 * 14a/3 = 7a/3 area = sqrt(s(s-a)(s-b)(s-c)) 8sqrt(14) = sqrt( 7a/3 * (7a/3 - a)(7a/3 - 5a/3)(7a/3 - 2a)) 8sqrt(14) = sqrt(7a/3 * 4a/3 * 2a/3 * a/3) from here, we can solve a first

OpenStudy (wolf1728):

You still think it is solvable?

OpenStudy (anonymous):

oh wait !!!!!!!!! we made a big mistake

OpenStudy (raden):

yes that works. i got a = 6, b = 10, and c = 12

OpenStudy (anonymous):

@RadEn is right, that's not the ratio of sides but a combination of sides :-))

OpenStudy (anonymous):

@wolf1728 your work is fine so :-)

OpenStudy (wolf1728):

RadEn the ratios are supposed to be 4: 2: 1 6 : 10: 12 doesn't work.

OpenStudy (anonymous):

oh no, problem is 6b+c:3b+a:2c-a=4:2:1 not a:b:c=4:2:1

OpenStudy (wolf1728):

okay - that was my problem from the beginning

OpenStudy (raden):

(6b+c) : (3b+a) : (2c-a) = 4 : 2 : 1 plug in the values of a=6, b=12, and c=12 the ratio be right :)

OpenStudy (raden):

i mean b=10

OpenStudy (anonymous):

for asker @someoneelse , see replies 1, 5 and 15

OpenStudy (wolf1728):

Hey we ALL got medals!

OpenStudy (anonymous):

:-))

OpenStudy (wolf1728):

Well I think I'll get going. See you later folks.